Make New Entry,
Make Followup Entry
User name A. Deur
Log entry time 16:01:20 on November21,2009
Entry number 301630
keyword=Day Shift summary
8:00 Start the shift with 99.7 uA in the Hall. Smooth production but Compton HV down. RHRS run #:5208, LHRS:25657, Parity:14112
8:18 runs completed
8:29 Start runs 5209, 25658, 14113
8:38 Noticed (thanks to the shift checklist) that the numbers on the Happex Feedback monitor were highlighted in red. Stopped and restarted the monitor and the problem cleared itself.
8:54 Stopped parity run earlier (44319 events) because we accidentally lost CODA for the HRSs (human mistake)
8:55 Start runs 5210, 25659, 14114
9:28 All CODA windows accidentally closed again. Restarting CODAs
9:32 Start runs 5211, 25660, 14115
10:02 Diancheng noticed that we are getting all the triggers for the above run instead of just T1. Maybe something was downlowad improperly or the prescales are not set correctly. Deadtime looks fine (4 and 7%) so these runs can be considered production.
10:28 Runs completed.
10:34 Start runs 5214, 25663, 14116
11:28 Start runs 5215, 25664, 14117
12:26 Start runs 5216, 25665, 14118
1:24 Start runs 5217, 25666, 14119
2:24 Start runs 5218, 25667, 14120
3:16 End runs early because a cryomodul is dow, PVDIS run has 93k evnets.
We are requesting an opportunistic controlled access for working on the scalers. Meanwhile we are taking pedestal runs, until DAQ work starts.